OpenStudy (anonymous):

What is the percentage markup on the cost of a bicycle if the retail price is $140 and the cost is $80? 7.5% 17.5% 1.75% .75<--my answer What is the percentage markup on the retail price of a pair of hiking boots if the retail price is $80 and the cost is $46? 4.25% .425 .360 6.30%

OpenStudy (anonymous):

@Loser66

OpenStudy (vikstar2.0):

1.75 for the first problem

OpenStudy (anonymous):

ty and any idea on the second one?

OpenStudy (vikstar2.0):

.425 is the answer to the second

OpenStudy (vikstar2.0):

so was i right? if i was can i get a medal?

OpenStudy (anonymous):

Using the cost method of pricing, what is the retail price of a music CD if the cost is $6.75 and the markup on cost is .35? Please round the answer to the nearest cent. $7.74<---my answer $10.55 $8.43 $9.11

OpenStudy (anonymous):

is my answer correct to this?

OpenStudy (anonymous):

@hartnn

OpenStudy (vikstar2.0):

i thinks its $9.11

hartnn (hartnn):

Cost * %markup + cost = retail price (140-80)/80 = 60/80 = 0.75 for the 1st (80-46)/46 = 0.74 hmmm..i get 0.74 for 2nd which is not in the choices. 6.75*0.35 + 6.75 = 9.11 for 3rd

OpenStudy (anonymous):

hmm so what should i do for the second xD

hartnn (hartnn):

try 0.425...